# Cayley Graph Propagation

### v1

@@ 6 Conclusion
In this work, we presented Cayley Graph Propagation (CGP), an efficient propagation scheme that mitigates over-squashing. CGP utilises the complete Cayley Graph structure to guarantee improved information flow between nodes in the input graph. We highlight the advantages of Cayley graphs as expanders and regular graphs. We show that by truncating the Cayley graphs to align with the input graph, as suggested in Expander Graph Propagation, the resulting graph may contain bottlenecks. This is in contrast to the Cayley graph we use, which is guaranteed to be bottleneck-free. We demonstrate the effectiveness and efficiency of CGP compared to EGP and other rewiring approaches, over multiple real-world datasets, including large-scale and long-range datasets.
@@
One limitation of our proposed model is the performance on datasets containing graphs with a comparatively higher node-to-edge ratio. For this reason, one such avenue for future work is aligning the Cayley graph edges such that that they retain the inductive bias of the task [33]. Additionally, it would be interesting to see how CGP performs in other tasks that utilise expander graphs, including but not limited to temporal graph rewiring [31]. Furthermore, concurrent work has used virtual nodes as the focal point of their proposed methods [53, 54] with Southern et al. [55] analysing virtual nodes’ role within the context of over-squashing. Therefore, we theorise an interesting setting would be applying these authors’ approaches to the additional virtual nodes retained from the complete Cayley graph structure.
Appendix A Experimental Details
@@
Table 5: Statistics of the datasets from OGB, TUDataset and LRGB used as part of our evaluation.
Dataset #Graphs Average #Nodes Average #Edges #Classes
OGBG-MOLHIV 41,127 25.5 27.5 2
OGBG-PPA 158,100 243.4 2,266.1 37
REDDIT-BINARY 2,000 429.6 995.5 2
IMDB-BINARY 1,000 19.8 193.1 2
MUTAG 188 17.9 39.6 2
ENZYMES 600 32.6 124.3 6
PROTEINS 1,113 39.1 145.6 2
COLLAB 5,000 74.5 4,914.4 3
PEPTIDES-FUNC 15,535 150.94 307.30 10
PEPTIDES-STRUCT 15,535 150.94 307.30 11
@@
In this section, we provide thorough details regarding our experimental setting for each of our datasets. We utilise the well-established experimental settings [44, 9] that are used for OGB and TUDataset respectively. The experimental setting for the LRGB [43] is found in Section B. Refer to Table 5 for the dataset statistics. Of significance, to rule out performance gain due to hyperparameter tuning in our results we use the same setting of GNN and hyperparameters for each task and model. All of our experiments use the default settings of the Adam optimiser [56] with a learning rate of $1\times 10^{-3}$, however the TUDataset and LRGB use the ReduceLROnPlateau scheduler with differing parameters.
A.1 Datasets
@@
From the OGB [44] we consider two of the graph classification tasks: ogbg-molhiv and ogbg-ppa. ogbg-molhiv is among the largest molecule property prediction datasets within the scope of the MoleculeNet benchmark [46], and ogbg-ppa focuses on classifying species based on their taxa, using their protein-protein association networks [47]. The OGB provides a unified evaluation protocol for each dataset, including application-specific data splits and evaluation metrics. Accordingly, our experimental setup to empirically evaluate against the OGB datasets leverages the official open-source implementation from the OGB authors [44]. ogbg-molhiv uses a 80%/10%/10% train/validation/test split and ROC-AUC for the evaluation metric, whilst ogbg-ppa uses a 50%/28%/22% train/validation/test split and accuracy for the evaluation metric. The hyperparameter setting for our models includes 5 layers, a hidden dimension of 300, a dropout of 50% and the use of Batch Norm [57]. In Table 1 the results reported for our model are trained to 100 epochs across 10 seeds and 5 seeds for ogbg-molhiv and ogbg-ppa respectively.
@@
The TUDataset [42] is considered under the claim of Karhadkar et al. [9] that the topology of the graphs in relation to the tasks require long-range interactions. Thus, we consider all graph classification tasks from the TUDataset: REDDIT-BINARY, IMDB-BINARY, MUTAG, ENZYMES, PROTEINS and COLLAB. Our experimental setup for the TUDataset is akin to the well-established and open-source setting of Karhadkar et al. [9]. Accordingly, we train our GNNs with 80%/10%/10% train/validation/test split and use a stopping patience of 100 epochs based on the validation loss. The ReduceLROnPlateu uses the default setting as found in PyTorch [58]. However unlike Karhadkar et al. [9], we report the accuracy over 20 random seeds and set the maximum number of epochs to 300 [48]. Additionally, in line with our OGB setup we apply Batch Norm [57]. The results reported in Table 2 use a model that is more akin to our OGB experimental setup, therefore our hyperparameters include setting the number of layers to 5, a hidden dimension of 128 and a dropout of 10%.
For Table 3, we use the hyperparameters as in the well-established graph rewiring baselines: DIGL [40], SDRF [8], FoSR [9], BORF [41] and GTR [11]. By adopting the hyperparameters of these baselines methods, we not only ensure a fair comparison, but demonstrate that CGP remains competitive even under their prescribed conditions. To this end, in Table 3 and as in Karhadkar et al. [9] we use 4 layers, a hidden dimension of 64 and a dropout of 50%. Additionally, in accordance with Karhadkar et al. [9] the results in Table 3 are reported to a 95% confidence interval. We report the standard deviation $\sigma$ for all other tables, due to this being more commonly used within the community. For the graph rewiring techniques, we follow the hyperparameters as reported in the respective baselines. This includes the teleport probability ($\alpha$) and sparsification threshold ($\epsilon$) for DIGL, as well as the number of rewiring iterations for SDRF and FoSR being derived from Karhadkar et al. [9]. For BORF this includes the number of batches ($n$), number of edges added per batch ($h$), and number of edges removed per batch ($k$) from Nguyen et al. [41]. Finally, for GTR this includes the number of edges added from Black et al. [11].
A.2 Hardware
@@
Appendix B Additional Experiments
@@ Results
Appendix C Scalability Analysis of CGP
@@
We empirically analyse the scalability of CGP by comparing the computational preprocessing time against the state-of-the-art graph rewiring techniques [40, 8, 9, 41, 11]. We first provide a real-world evaluation by benchmarking the preprocessing time on two real-world datasets: TUDataset [42] and LRGB [43]. The results are reported in Table 7 and 8, using the same graph rewiring techniques as in Table 3 and 6 respectively. To extend our scalability analysis, we create a synthetic benchmark by leveraging Erdős–Rényi with a probability $p=\frac{5\log n}{n}$ (as used by Karhadkar et al. [9]) to create graphs of up to 10,000 nodes. In line with the results reported in Table 3, we do not conduct the synthetic benchmark for BORF [41], due to the impracticality of the rewiring time which is highlighted in Section 5.
@@
The results show the efficacy of our proposed CGP model, as the preprocessing time is orders of magnitude lower than the graph rewiring techniques. To this end, we examine the lack of computational preprocessing time required to generate the corresponding Cayley graphs for both CGP and EGP. Overall, our results show the practicality of CGP to scale to large graphs when compared to the graph rewiring approaches. This is reinforced by the experimentation being conducted on the local machine with leading hardware as in Section A.2. In particular, both the CPU and GPU deliver top-tier clock speeds, therefore on lower-performing hardware, the graph rewiring techniques could have a more detrimental effect.
Appendix D Effective Resistance of Cayley graphs
(a) MUTAG
(b) ENZYMES
(c) PROTEINS
(d) IMDB-BINARY
(e) COLLAB
(f) OGBG-MOLHIV
The results in Figure 5 show the average of the total effective resistance $R_{tot}$ for all the corresponding Cayley graphs against the base input graphs and truncated Cayley graphs as found in EGP. Akin to the results presented in Black et al. [11], for a fair evaluation we do not include graphs that may be disconnected. This is because the Cayley graphs used in CGP are a complete and regular graph structure, therefore every node will be connected. The results show that CGP consistently has a lower total effective resistance $R_{tot}$ in comparison to EGP. Significantly, the complete Cayley graph structure is chosen by recalling $|V(\mathrm{Cay}(\mathrm{SL}(2,\mathbb{Z}_{n});S_{n}))|\geq|V|$, therefore the total effective resistance $R_{tot}$ for CGP may be inflated due to it potentially being summed over more pairs of nodes. This further reinforces our claim that it is more beneficial to use the complete Cayley graph structure with the additional nodes serving as shortcuts for message passing between nodes along the graph.
Appendix E Dirichlet Energy of Cayley graphs
@@
Here, we evaluate the impact of propagating over the complete Cayley graph structure in regards to the over-smoothing problem using the Dirichlet energy. As stated in Section 1, a trade-off occurs between mitigating over-squashing and causing over-smoothing. The Dirichlet energy quantifies over-smoothing by measuring the deviation of a function on the graph from being constant between connected node pairs, thus indicating the level of non-smoothness in the signals [34]. In turn, the Dirichlet energy has been used to measure the amount of over-smoothing in graph representations [9, 12, 48].
@@
Appendix F Additional Experiment Details for Section 3.2
Sum Task
The teacher readout is sampled once from $\mathcal{N}(0,1)$ and used for all the graphs. All graphs have $n=20$ nodes, and each node is assigned with a feature vector in $\mathbb{R}^{128}$ sampled i.i.d from $\mathcal{N}(0,1)$.
We used a $1$-layer “student" GNN following the teacher model, with readout and ReLU activations.
Appendix G Proof of Lemma 3.1
We consider the case of a node classification task over Cayley graphs. We wish to show that if the training data consists of infinitely many samples from a distribution over $\mathrm{Cay}(\mathrm{SL}(k,\mathbb{Z}_{n});S_{n})$, then the learned model will extrapolate perfectly to a distribution over graphs drawn from $\mathrm{Cay}(\mathrm{SL}(k,\mathbb{Z}_{n^{\prime}});S_{n^{\prime}})$, where $n\neq n$. We assume the same feature distribution in all cases.
Let $f$ be a node classifier with perfect accuracy on the support of the training distribution.
Let $G=(X_{G},A_{\mathrm{Cay}(\mathrm{SL}(k,\mathbb{Z}_{n});S_{n})})$ be in the support of the training distribution and $v$ a vertex in $G$. Let $G^{\prime}=(X_{G},A_{\mathrm{Cay}(\mathrm{SL}(k,\mathbb{Z}_{n^{\prime}});S_{n^{\prime}})})$ and $v^{\prime}$ a node in $G^{\prime}$.
To show that $f$ classifies we utilise Theorem 4.1 in  Yehudai et al. [38], which uses the notion of d-patterns. For the sake of completeness we include the definition of d-patterns and the theorem here.
Definition G.1.
Let C be a finite set of node features, and let $G=(V,E)$ be a graph with node feature $c_{v}\in C$ for every node $v\in V$ We define the d-pattern of a node $v\in V$ for $d\geq 0$ recursively: For $d=0$, the 0-pattern is $c_{v}$. For $d\geq 0$, the d-pattern of $v$ is $p=(p_{v},\{(p_{i_{1}},m_{p_{i_{1}}}),\dotsc,(p_{i_{l}},m_{p_{i_{l}}})\})$ iff node $v$ has $(d-1)$- pattern $p_{v}$ and for every $j\in\{1\dotsc l\}$ the number of neighbors of $v$ with $(d-1)$-pattern $p_{i_{j}}$ is exactly $m_{p_{i_{j}}}$. Here, $l$ is the number of distinct neighboring $d-1$ patterns of $v$.
Theorem G.2 (Theorem 4.2 from  Yehudai et al. [38]).
Any function that can be represented by a d-layer GNN is constant on nodes with the same d-patterns.
The regularity degree of nodes in graphs from $\mathrm{Cay}(\mathrm{SL}(k,\mathbb{Z}_{n^{\prime}})$ is the same as in nodes in graphs from $\mathrm{Cay}(\mathrm{SL}(k,\mathbb{Z}_{n});S_{n})$. Therefore, the d-patterns are the same. ∎

### v2

@@ 6 Conclusion
In this work, we presented Cayley Graph Propagation (CGP), an efficient propagation scheme that mitigates over-squashing. CGP utilises the complete Cayley Graph structure to guarantee improved information flow between nodes in the input graph. We highlight the advantageous topological properties of Cayley graphs for message-passing. We show that by truncating the Cayley graphs to align with the input graph, as suggested in Expander Graph Propagation, the resulting graph may contain bottlenecks. This is in contrast to the Cayley graph we use, which is guaranteed to be bottleneck-free. We demonstrate the effectiveness and efficiency of CGP compared to EGP and other rewiring approaches, over multiple real-world datasets, including large-scale and long-range datasets.
@@
One limitation of our proposed model is the performance on datasets containing graphs with a comparatively higher node-to-edge ratio. For this reason, one such avenue for future work is aligning the Cayley graph edges such that that they retain the inductive bias of the task [47]. Additionally, it would be interesting to see how CGP performs in other tasks that utilise expander graphs, including but not limited to temporal graph rewiring [25]. Furthermore, concurrent work has used virtual nodes as the focal point of their proposed methods [48, 49] with Southern et al. [31] analysing the role of virtual nodes within the context of over-squashing. Therefore, we theorise an interesting setting would be applying these authors’ approaches to the additional virtual nodes retained from the complete Cayley graph structure.
Acknowledgements
We would like to thank Petar Veličković’s co-authors of Expander Graph Propagation, including Andreea Deac and Mark Lackenby. Their novel work of leveraging the use of expander graphs within the context of graph neural networks to alleviate bottlenecks and counter over-squashing provided the basis for our work. Furthermore, we would like to thank Alex Vitvitskyi, Csaba Szepesvári, Johannes Vallikivi and Dobrik Georgiev for their invaluable insights prior to submission.
Appendix A Effective Resistance of Cayley graphs
(a) MUTAG
(b) ENZYMES
(c) PROTEINS
(d) IMDB-BINARY
(e) COLLAB
(f) OGBG-MOLHIV
The results in Figure 3 show the average of the total effective resistance $R_{tot}$ for all the corresponding Cayley graphs against the base input graphs and truncated Cayley graphs as found in EGP. Akin to the results presented in Black et al. [12], for a fair evaluation we do not include graphs that may be disconnected. This is because the Cayley graphs used in CGP are a complete graph structure, therefore every node will be connected. The results show that CGP consistently has a lower total effective resistance $R_{tot}$ in comparison to EGP. Significantly, the complete Cayley graph structure is chosen by recalling $|V(\mathrm{Cay}(\mathrm{SL}(2,\mathbb{Z}_{n});S_{n}))|\geq|V|$, therefore the total effective resistance $R_{tot}$ for CGP may be inflated due to it potentially being summed over more pairs of nodes. This further reinforces our claim that it is more beneficial to use the complete Cayley graph structure with the additional nodes serving as shortcuts for message passing between nodes along the graph.
Appendix B Virtual Nodes Initialisation
Table 5: Results of CGP using different virtual node initialisation strategies, including ZEROS, ONES, RANDOM and $\theta$ on the TUDataset. ZEROS is the baseline CGP and $\theta$ denotes the nodes are initialised using a learnable parameter.
Model REDDIT-BINARY IMDB-BINARY MUTAG ENZYMES PROTEINS COLLAB
+ CGP-ZEROS $78.225\pm 1.268$ $\textbf{71.650}\pm 1.532$ $\textbf{85.250}\pm 3.200$ $\textbf{50.083}\pm 2.242$ $\textbf{73.080}\pm 1.396$ $\textbf{73.350}\pm 0.788$
+ CGP-ONES $79.175\pm 1.992$ $71.000\pm 1.806$ $82.250\pm 3.773$ $47.833\pm 2.428$ $70.535\pm 1.972$ $73.200\pm 0.964$
+ CGP-RAND $\textbf{82.300}\pm 1.368$ $71.200\pm 1.532$ $84.000\pm 3.130$ $44.083\pm 2.891$ $70.000\pm 1.937$ $73.020\pm 0.742$
+ CGP-$\theta$ $80.150\pm 1.888$ $69.250\pm 2.034$ $85.000\pm 3.606$ $49.750\pm 2.752$ $70.133\pm 2.053$ $73.410\pm 0.977$
(a) MUTAG
(b) ENZYMES
(c) PROTEINS
Figure 4: The mean norm of the virtual node embeddings for CGP using different initialisation strategies on the TUDataset, including ZEROS, ONES, RANDOM and $\theta$.
(a) MUTAG
(b) ENZYMES
(c) PROTEINS
Figure 5: The variance in the norm of the virtual node embeddings for CGP using different initialisation strategies on the TUDataset, including ZEROS, ONES, RANDOM and $\theta$.
In Section 4, we show that virtual nodes in our proposed CGP method provide the flexibility to be initialised in some pre-defined manner. For the baseline CGP model, we opted to initialise them to zeros, which is in line with the work of [5, 30, 31]. In this section, we empirically evaluate whether different node initialisation strategies impact the performance of CGP. We report the results in Table 5, leveraging the TUDataset with the hyperparameter setting as prescribed in Section C. For our other node initialisation strategies, we choose constant ones, random features that are sampled from $\mathcal{N}(0,1)$ and $\theta$ where $\theta$ is a learnable parameter. Notably, ones is an interesting initialisation strategy as for datasets with no node features they are commonly assigned to a constant of ones [36]. The results overall show that our baseline model of CGP-ZERO performs the best across all datasets, however all approaches obtain similar performance within the variance of each other.
Next, we investigate the role of the additional virtual nodes in CGP. To this end, we use the mean norm of the virtual node embeddings, as well as the variance between them as a statistical proxy to measure the diversity of the embeddings. We visualise the results in Figure 4 and 5 respectively for all of the virtual node initialisation strategies found in Table 5, however we fix the number of epochs to $100$ and report the average over 10 random seeds. The results show that during the training lifetime of the model, the mean norm and variance of virtual node node embeddings grow, indicating that they are learning distinct representations. Overall, the results in this section show that the initialisation of the virtual nodes impact CGP. However, due to the focus of CGP being to use complete Cayley graph structure and the additional virtual nodes serving as a conduit to this end. Therefore, a deeper understanding of the role of the sparsely connected virtual nodes and their initialisation strategies presents an interesting avenue for future work.
Appendix C Experimental Details
@@
Table 6: Statistics of the datasets from OGB, TUDataset and LRGB used as part of our emperical evaluation. denotes the number of additional virtual nodes (VN) added in our CGP model.
Dataset #Graphs Average #Nodes Average #Edges #Classes Average #VN Max #VN
OGBG-MOLHIV 41,127 25.5 27.5 2 11.8 191
OGBG-PPA 158,100 243.4 2,266.1 37 50.3 191
REDDIT-BINARY 2,000 429.6 995.5 2 139.4 1809
IMDB-BINARY 1,000 19.8 193.1 2 10.7 71
MUTAG 188 17.9 39.6 2 8.0 23
ENZYMES 600 32.6 124.3 6 14.6 71
PROTEINS 1,113 39.1 145.6 2 20.6 190
COLLAB 5,000 74.5 4,914.4 3 34.2 260
PEPTIDES-FUNC 15,535 150.94 307.3 10 67.0 263
PEPTIDES-STRUCT 15,535 150.94 307.3 11 67.0 263
@@
In this section, we provide thorough details regarding our experimental setting for each of our datasets. We utilise the well-established experimental settings [38, 10] that are used for OGB and TUDataset respectively. The experimental setting for the LRGB [37] is found in Section D. Refer to Table 6 for the dataset statistics. Of significance, to rule out performance gain due to hyperparameter tuning in our results we use the same setting of GNN and hyperparameters for each task and model. All of our experiments use the default settings of the Adam optimiser [51] with a learning rate of $1\times 10^{-3}$, however the TUDataset and LRGB use the ReduceLROnPlateau scheduler with differing parameters.
C.1 Datasets
@@
From the OGB [38] we consider two of the graph classification tasks: ogbg-molhiv and ogbg-ppa. ogbg-molhiv is among the largest molecule property prediction datasets within the scope of the MoleculeNet benchmark [40], and ogbg-ppa focuses on classifying species based on their taxa, using their protein-protein association networks [41]. The OGB provides a unified evaluation protocol for each dataset, including application-specific data splits and evaluation metrics. Accordingly, our experimental setup to empirically evaluate against the OGB datasets leverages the official open-source implementation from the OGB authors [38]. ogbg-molhiv uses a 80%/10%/10% train/validation/test split and ROC-AUC for the evaluation metric, whilst ogbg-ppa uses a 50%/28%/22% train/validation/test split and accuracy for the evaluation metric. The hyperparameter setting for our models includes 5 layers, a hidden dimension of 300, a dropout of 50% and the use of Batch Norm [42]. In Table 1, the results reported for our model are trained to 100 epochs across 10 seeds and 5 seeds for ogbg-molhiv and ogbg-ppa respectively.
@@
The TUDataset [36] is considered under the claim of Karhadkar et al. [10] that the topology of the graphs in relation to the tasks require long-range interactions. Thus, we consider all graph classification tasks from the TUDataset: REDDIT-BINARY, IMDB-BINARY, MUTAG, ENZYMES, PROTEINS and COLLAB. Our setup for all TUDataset experiments is akin to the well-established and open-source setting of Karhadkar et al. [10]. Accordingly, we train our GNNs with 80%/10%/10% train/validation/test split and use a stopping patience of 100 epochs based on the validation loss. We fix the number of layers to 4 with a hidden dimension of 64 and a dropout of 50%. The ReduceLROnPlateu uses the default setting as found in PyTorch [52]. Additionally, in accordance with Karhadkar et al. [10] the results for the TUDataset are reported to a 95% confidence interval. However unlike Karhadkar et al. [10], we report the accuracy over 20 random seeds, set the maximum number of epochs to 300 [35], as well as apply Batch Norm [42].
For TUDataset experiments, we compare CGP against the state-of-the-art approaches that require dedicated preprocessing and use the hyperparameters as in the well-established baselines: DIGL [33], SDRF [9], FoSR [10], BORF [34], GTR [12] and PANDA [35]. By adopting the hyperparameters of these baselines methods, we not only ensure a fair comparison, but demonstrate that CGP remains competitive even under their optimised settings. For the graph rewiring techniques, we follow the hyperparameters as reported in the respective baselines. This includes the teleport probability ($\alpha$) and sparsification threshold ($\epsilon$) for DIGL, as well as the number of rewiring iterations for SDRF and FoSR being derived from Karhadkar et al. [10]. For BORF this includes the number of batches ($n$), number of edges added per batch ($h$), and number of edges removed per batch ($k$) from Nguyen et al. [34]. For GTR this includes the number of edges added from Black et al. [12]. Finally, this includes the hyperparameters for PANDA from Choi et al. [35]: the centrality metric $C(G)$, the top-$k$ nodes and the increased width (i.e. hidden dimension) of the model.
C.2 Hardware
@@
Appendix D Additional Experiments
@@ Results
Appendix E Scalability Analysis of CGP
@@
We empirically analyse the scalability of CGP by comparing the computational preprocessing time against the state-of-the-art graph rewiring techniques [33, 9, 10, 34, 12], as well as PANDA [35]. We first provide a real-world evaluation by benchmarking the preprocessing time on two real-world datasets: TUDataset [36] and LRGB [37]. The results are reported in Table 8 and 9, using the same graph rewiring techniques as in Table 2 and 7 respectively. To extend our scalability analysis, we create a synthetic benchmark by leveraging Erdős–Rényi with a probability $p=\frac{5\log n}{n}$ (as used by Karhadkar et al. [10]) to create graphs of up to 10,000 nodes. In line with the results reported in Table 2, we do not conduct the synthetic benchmark for BORF [34], due to the impracticality of the rewiring time which is highlighted in Section 5.
@@
The results show the efficacy of our proposed CGP model, as the preprocessing time is orders of magnitude lower than the graph rewiring techniques. To this end, we examine the lack of computational preprocessing time required to generate the corresponding Cayley graphs for both CGP and EGP. Overall, our results show the practicality of CGP to scale to large graphs when compared to the graph rewiring approaches. This is reinforced by the experimentation being conducted on the local machine with leading hardware as in Section C.2. In particular, both the CPU and GPU deliver top-tier clock speeds, therefore on lower-performing hardware, the graph rewiring techniques could have a more detrimental effect.
Appendix F Cayley graphs as Regular graphs
In this section, we explore the additional benefits of Cayley graphs being a regular graph. It was observed by Bechler-Speicher et al. [26] that GNNs tend to overfit the given graph structure, even in cases where it does no provide useful information for the predictive task. Nonetheless, it was shown that regular graphs exhibit robustness to this overfitting. As Cayley graphs are regular graphs, they exhibit this robustness.
We used four different datasets from this baseline by sampling graph-structures from different graph distributions. The set of node feature vectors remains the same across all the datasets, and thus, the datasets differ only in their graph structure. The graph distributions we used are: Cayley graphs over $24$ nodes, star-graph (Star) where the only connections are between one specific node and all other nodes, and the preferential attachment model (BA)  [56], where the graph is built by incrementally adding new nodes and connecting them to existing nodes with probability proportional to the degrees of the existing nodes. We used the data as is with empty graphs (Empty) as a baseline to compare to. On each dataset, we varied the training set size and evaluated test errors on $5$ runs with random seeds.
Extrapolation
The ability and failures of GNNs to extrapolate to graphs of sizes larger then the one presented during training was examined in Yehudai et al. [57]. It was shown that size generalisation is dependent on local structures around each node, called d-patterns. In particular, if increasing the graph size does not change the distribution of these d-patterns, then extrapolation to larger graph sizes is guaranteed.
Sum Task
The teacher readout is sampled once from $\mathcal{N}(0,1)$ and used for all the graphs. All graphs have $n=20$ nodes, and each node is assigned with a feature vector in $\mathbb{R}^{128}$ sampled i.i.d from $\mathcal{N}(0,1)$. We used a $1$-layer “student" GNN following the teacher model, with readout and ReLU activations.
Appendix G Dirichlet Energy of Cayley graphs
@@
Here, we evaluate the impact of propagating over the complete Cayley graph structure in regards to the over-smoothing problem using the Dirichlet energy. Over-smoothing is an independent problem from over-squashing, but another well-known problem that impacts the expressivity of GNNs [59, 60]. This phenomenon occurs in GNNs when the number of layers increases [61, 62], such that node features become increasingly similar [63]. However, the over-smoothing problem is linked with over-squashing due to a common approach to the latter being graph rewiring; too many additional edges lead to over-smoothing [10]. There are varying approaches to measure over-smoothing for a graph with one such notable metric being the Dirichlet energy [64, 10, 13].
The Dirichlet energy quantifies over-smoothing by measuring the deviation of a function on the graph from being constant between connected node pairs, thus indicating the level of non-smoothness in the signals [27]. In turn, the Dirichlet energy has been used to measure the amount of over-smoothing in graph representations [10, 13, 35].